OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 80°, a = 24, b = 50

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

That one and Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, 2>, v = <21, 6> if you know how to do um >.<

OpenStudy (unklerhaukus):

The law of sines say that for a triangle with angles A,B,C and with sides a,b,c \[\frac{\sin A}{a}=\frac{\sin B}{b}=\frac{\sin C}{c}\]